OpenStudy (anonymous):

Anthony leaves his house on his bike traveling at 18 mph. Thirty minutes later, his brother Chandler takes after him at 24 mph. How many hours will it take Chandler to catch up to Anthony?

OpenStudy (anonymous):

I've got d=18*30+24*30 540+720 and my answer I got was 1260

OpenStudy (anonymous):

distance = rate * time if Chandler is going to catch up to Anthony, thanthis means they will travel the same distance. so, distance anthony = distance chandler rate * time of anthony = rate * time of chandler anthony leaves 30 minutes before chandler, so we do t+30 for anthony's time 18*(t+30) = 24*t

OpenStudy (jdoe0001):

\(\begin{array}{llll} &distance&rate&time\\ \hline\\ anthony&&18&h\\ chandler&&24&h-30 \end{array}\) by the time Chandler starts going, Anthony has already been traveling for 30mins, so Chandler's time will be whatever Anthony's is, behind him by 30mins they both meet when the distances are equal to each other

OpenStudy (anonymous):

Im still confused

OpenStudy (anonymous):

does this idea make sense? 'if Chandler is going to catch up to Anthony, thanthis means they will travel the same distance. so, distance anthony = distance chandler'

OpenStudy (jdoe0001):

say your sister goes to school at 8am, she forgets her protractor so at 8:30am, you find out she forget her protractor, so you go after her you catch up with her at 8:45am, she's been walking from 8:00am to 8:45am, so 45mins you've been running from 8:30am to 8:45am, so 15mins you catch up with her 2 miles from home, that means she walked 2 miles and you walked 2 miles the distances must match, thus the meaning of catch up